# Multistatic Radar Performance in the Presence of Distributed Wireless Synchronization

### v2

@@ I Introduction
Multistatic radars [11, 17] (MSRs) offer significant advantages in detecting stealth targets that are designed to scatter energy away from the monostatic radar line of sight (LOS). Passive mode operation makes them inherently more covert and resilient to electronic countermeasures. The growing deployment of unmanned aerial vehicles (UAVs) and vehicular nodes further enhances the appeal of MSR architectures for modern sensing and surveillance applications. However, their performance critically depends on achieving precise synchronization and maintaining accurate knowledge of the cooperating radar locations. Prior research has explored various synchronization strategies, including wired [13], GPS/GNSS-based [22, 21, 3], and GPS/GNSS-independent wireless approaches [2, 15, 16, 23].
@@
In GPS/GNSS-based approaches, the local oscillator (LO) that drives the digital subsystems, such as ADCs, DACs, carrier-generation LOs, and data clocks, is disciplined by a GPS/GNSS timing receiver, commonly referred to as a GPS-disciplined oscillator (GPSDO). These units provide stable timing and frequency references, with typical timing accuracies of 5–20 ns and fractional frequency stability on the order of $10^{-13}$$10^{-11}$ [3, 14], achieving long-term stability. The short-term stability is determined by the oscillator used and its phase-noise characteristics. However, GPS signals can be easily spoofed or jammed, require LOS visibility to satellites, and typically need a long acquisition time to achieve lock. Moreover, high-performance GPSDOs are expensive, although low-cost commercial off-the-shelf (COTS) options are available with moderate stability.
@@
To overcome the dependency on GPS/GNSS, independent wireless synchronization approaches have been developed, which can be broadly classified into centralized [2, 23, 15, 16] and decentralized [24] architectures. These are designed for general distributed wireless communication systems, with a focus on improving synchronization performance and demonstrated sub-hertz frequency accuracy and nanosecond-to-picosecond (ns–ps) time accuracy. In centralized systems, a leader–follower architecture is adopted, where the followers synchronize to the leader by exchanging frequency and time reference signals. Frequency synchronization techniques typically rely on the wireless transmission of an analog [1, 2, 15, 16, 5] or digitally modulated reference from the leader. Time synchronization, on the other hand, is achieved through bi-directional waveform (BDW) exhanges [2, 23, 15, 16, 4, 6, 18], which can be implemented using either timestamp-based or timestamp-free protocols. A key challenge for frequency synchronization is its susceptibility to multi-path and Doppler effects, while time synchronization schemes require symmetric propagation delays between the leader and followers. Furthermore, the achievable accuracy is strongly dependent on the SNR of the wireless link.
@@
Prior studies have investigated GPS-based synchronization for MSR from different perspectives. Coherent network radar operation using custom GPS-disciplined oscillators was demonstrated in [22]. The follow-up work [21] analyzed synchronization limitations and proposed line-of-sight phase compensation for bistatic coherence. The work in  [3] evaluated commercial GNSS-disciplined oscillators for practical radar synchronization. Wireless adaptation of the white rabbit precision time protocol (WR-PTP) was studied in [19, 8], with a particular focus on synchronization. In contrast, this paper investigates the performance of a leader–follower MSR architecture employing GPS/GNSS-independent wireless synchronization. A Bayesian Cramér–Rao lower bound (BCRLB) [20, 25, 26] framework is developed to study the impact of the residual synchronization offsets on target position and velocity estimation accuracy. In this work, frequency synchronization is achieved by extracting a reference clock from a two-tone waveform (TTW) transmitted by leader to all followers [1, 2, 15, 16, 5]. Time synchronization is accomplished using a timestamp-free bidirectional waveform exchange that estimates and corrects clock offsets while compensating for propagation delays [4, 6, 18]. Simulation results demonstrate the MSR performance under multiple configurations of synchronization-link channel and transmit parameters, and compared with the monostatic radar.
@@
The paper is structured as follows. First, the system model incorporating synchronization offsets is introduced. Next, the BCRLB framework is presented, where the synchronization offsets are treated as nuisance parameters with prior distributions. The subsequent section describes the MSR geometry and derives the corresponding CRLBs for target position and velocity. The procedures for frequency and time synchronization are then outlined. Finally, the simulation results are discussed, emphasizing the synchronization integrated MSR’s performance, followed by the conclusion and furture works.
@@ II Signal Model with Synchronization Offsets
In this section, we present the signal model incorporating synchronization offsets along with the underlying assumptions used in this paper. Fig. 1 illustrates a MSR architecture featuring inter-radar wireless links for synchronization and data exchange, as well as radar links between the radars and the target. The followers communicate wirelessly with the leader to achieve synchronization and to exchange Time of Arrival (TOA) and Frequency of Arrival (FOA) measurements. The radars performing TOA and FOA measurements are assumed to be either mobile or fixed wireless nodes with known locations relative to the leader. The focus of this work is on the impact of synchronization errors, while the position uncertainty of the followers is neglected. The followers, considered passive radars (e.g., drones or vehicular nodes), are geographically distributed around the leader, providing scalability in the number of nodes and adaptability to mobility based on operational requirements.
The transmitted signal from the leader radar is expressed as $s_{p}(t)=s(t)e^{j2\pi f_{c}t},$ where $s(t)$ denotes the complex baseband radar pulse defined over the interval $t\in\left[-\frac{T}{2},\frac{T}{2}\right]$ with a duration of $T$ seconds, and $f_{c}$ is the carrier frequency. Assuming a LOS channel between the transmitter and target as well as receiver and target and neglecting synchronization offsets, the received signal at the $i$th follower after down-conversion can be written as
@@
In practical systems, synchronization is typically performed parallel or in advance to target detection, to make sure that the carrier frequency offsets (CFOs) and clock (time) offsets (TOs) are corrected, ensuring carrier and timing alignment among all cooperating radars. However, residual carrier frequency offsets and clock offsets may still remain, resulting in frequency and timing mismatches across the radars. Incorporating these residual offsets into the ideal received signal and accounting for complex additive white Gaussian noise (AWGN), $n_{i}(t)\sim\mathcal{CN}(0,N_{0})$, the received signal after down-conversion at the $i$th radar can be expressed as
@@
where $\delta t_{i}$ and $\delta f_{i}$ represent the residual time and frequency offsets, respectively and will be treated as nuisance parameters [9]. We combine the amplitude term $\sqrt{G_{i}}$ and the constant phasor $h_{i}=h_{i}^{r}+jh_{i}^{i}=e^{j(-2\pi f_{i}\tau_{i}+\theta_{i})}$ into a single complex channel coefficient $A_{i}\triangleq\sqrt{G_{i}}h_{i}$. The received signal can thus be expressed as
@@
The terms $G_{i}$ and $h_{i}$ are included in the effective SNR, and since $|h_{i}|^{2}=1$, the received SNR is given by $\gamma_{i}=2G_{i}\frac{E_{s}}{N_{0}}$. The deterministic parameters of interest are the TOA, $\tau_{i}$, and the FOA, $f_{i}$. The residual clock and frequency offsets are modeled as independent Gaussian random variables, $\delta t_{i}\sim\mathcal{N}(0,\sigma_{t_{i}}^{2})$ and $\delta f_{i}\sim\mathcal{N}(0,\sigma_{f_{i}}^{2})$, with variances determined by the performance of the employed wireless synchronization techniques. The deterministic mean of the received (noise-free) signal is then given by $\mu_{i}(t)\triangleq\mathbb{E}\{r_{i}(t)\}=A_{i}\,s\!\left(t-(\tau_{i}+\delta t_{i})\right)e^{j2\pi(f_{i}+\delta f_{i})t}$.
@@ III BCRLB for Delay $\tau$ and Doppler $f_{d}$
We present the lower bounds on the variances of delay and Doppler estimation in the presence of synchronization offsets. The BCLRB [20, 25, 26] provides a lower bound on the mean squared error (MSE) of any estimator, whether unbiased or biased, when prior information about the unknown parameters is available. It extends the classical CRLB by incorporating information from both the likelihood function and the parameter’s prior distribution. Let $\boldsymbol{x}\in\mathcal{X}\subset\mathbb{R}^{D}$ denote a random parameter vector of interest, and let $\boldsymbol{y}\in\mathcal{Y}\subset\mathbb{R}^{K}$ represent the corresponding observation vector obtained from a statistical model parameterized by $\boldsymbol{x}$. Let $\hat{\boldsymbol{x}}(\boldsymbol{y})$ denote an estimator of $\boldsymbol{x}$. The BCRLB is defined as
$\mathbb{E}_{\boldsymbol{x},\boldsymbol{y}}\!\left[(\hat{\boldsymbol{x}}(\boldsymbol{y})-\boldsymbol{x})(\hat{\boldsymbol{x}}(\boldsymbol{y})-\boldsymbol{x})^{\top}\right]\succeq\mathbf{V}_{B}\triangleq\mathbf{J}_{B}^{-1},$ (4)
where $\mathbf{V}_{B}\in\mathbb{R}^{D\times D}$ denotes the BCRB matrix, and $\mathbf{J}_{B}\in\mathbb{R}^{D\times D}$ is the Bayesian information matrix (BIM), defined as
$\mathbf{J}_{B}\triangleq\mathbb{E}_{\boldsymbol{x},\boldsymbol{y}}\!\left[\nabla_{\boldsymbol{x}}\log p(\boldsymbol{x},\boldsymbol{y})\,\nabla_{\boldsymbol{x}}^{\top}\log p(\boldsymbol{x},\boldsymbol{y})\right].$ (5)
The BIM can be decomposed into two components, $\mathbf{J}_{B}=\mathbf{J}_{P}+\mathbf{J}_{D},$ where $\mathbf{J}_{P}$ represents the prior information term, given by
$\mathbf{J}_{P}\triangleq\mathbb{E}_{\boldsymbol{x}}\!\left[\nabla_{\boldsymbol{x}}\log p(\boldsymbol{x})\,\nabla_{\boldsymbol{x}}^{\top}\log p(\boldsymbol{x})\right],$ (6)
and $\mathbf{J}_{D}$ denotes the average Fisher information matrix (FIM) associated with the observations,
$\mathbf{J}_{D}\triangleq\mathbb{E}_{\boldsymbol{x},\boldsymbol{y}}\!\left[\nabla_{\boldsymbol{x}}\log p(\boldsymbol{y}|\boldsymbol{x})\,\nabla_{\boldsymbol{x}}^{\top}\log p(\boldsymbol{y}|\boldsymbol{x})\right]=\mathbb{E}_{\boldsymbol{x}}\!\left[\mathbf{J}_{F}(\boldsymbol{x})\right],$ (7)
with FIM as $\mathbf{J}_{F}(\boldsymbol{x})\triangleq\mathbb{E}_{\boldsymbol{y}|\boldsymbol{x}}\!\left[\nabla_{\boldsymbol{x}}\log p(\boldsymbol{y}|\boldsymbol{x})\,\nabla_{\boldsymbol{x}}^{\top}\log p(\boldsymbol{y}|\boldsymbol{x})\right].$
@@ III-A1 Fisher Information
The FIM for the parameter vector $\boldsymbol{x}=[\tau_{i},f_{i},\delta t_{i},\delta f_{i}]$ is defined as
$\mathbf{J}_{F}^{i}(\boldsymbol{x})=\begin{bmatrix}I_{\tau_{i}\tau_{i}}&I_{\tau_{i}f_{i}}&I_{\tau_{i}\delta t_{i}}&I_{\tau_{i}\delta f_{i}}\\ I_{f_{i}\tau_{i}}&I_{f_{i}f_{i}}&I_{f_{i}\delta t_{i}}&I_{f_{i}\delta f_{i}}\\ I_{\delta t_{i}\tau_{i}}&I_{\delta t_{i}f_{i}}&I_{\delta t_{i}\delta t_{i}}&I_{\delta t_{i}\delta f_{i}}\\ I_{\delta f_{i}\tau_{i}}&I_{\delta f_{i}f_{i}}&I_{\delta f_{i}\delta t_{i}}&I_{\delta f_{i}\delta f_{i}}\end{bmatrix}$.
@@ III-A3 Full FIM with Nuisance Parameters
Since $\frac{\partial\mu_{i}}{\partial\tau_{i}}=\frac{\partial\mu_{i}}{\partial\delta t_{i}}$ and $\frac{\partial\mu_{i}}{\partial f_{i}}=\frac{\partial\mu_{i}}{\partial\delta f_{i}}$, the full FIM can be written as $\mathbf{J}_{F}^{i}=\begin{bmatrix}\mathbf{A}^{i}&\mathbf{A}^{i}\\ \mathbf{A}^{i}&\mathbf{A}^{i}\end{bmatrix}$, where $\mathbf{A}^{i}$ denotes the FIM corresponding to the delay–Doppler parameters. In this work, the transmitted radar waveform is assumed to be a sequence of linear frequency modulated (LFM) pulses. The delay–Doppler FIM expressions derived in [7] are used directly to avoid re-derivation. The FIM is given by
$\mathbf{A}^{i}=2P\mathrm{SNR}_{1}\begin{bmatrix}\frac{1}{3}\pi^{2}f_{B}^{2}&-\frac{1}{6}\pi f_{B}T_{0}\\ -\frac{1}{6}\pi f_{B}T_{0}&\frac{1}{12}T_{0}^{2}\left(1+\left(\frac{T_{R}}{T_{0}}\right)^{2}(P^{2}-1)\right)\end{bmatrix}$ (9)
where $\mathrm{SNR}_{1}=\gamma_{i}\,\mathbf{a}(\theta)^{*}\mathbf{N}_{\Sigma}^{-1}\mathbf{a}(\theta)$ is the (post-processed [15]) SNR for a single pulse. Here, $P$ denotes the number of pulses, $T_{0}$ is the single-pulse duration, $f_{B}$ is the bandwidth of the chirp pulse (in the continuous domain), $T_{R}$ is the pulse repetition interval, $\mathbf{a}(\theta)$ is the array response vector, and $\mathbf{N}_{\Sigma}$ represents the spatial noise covariance matrix. Here, $\mathbf{J}_{D}^{i}=\mathbf{J}_{F}^{i}$, as $\mathbf{J}_{F}^{i}$ does not depend on the synchronization offsets.

### v3

@@ I Introduction
Multistatic radars [11, 16] (MSRs) offer significant advantages in detecting stealth targets that are designed to scatter energy away from the monostatic radar line-of-sight (LOS). Passive mode operation makes them inherently more covert and resilient to electronic countermeasures. The growing deployment of unmanned aerial vehicles (UAVs) and vehicular nodes further enhances the appeal of MSR for modern sensing and surveillance applications. However, their performance critically depends on achieving precise synchronization and maintaining accurate knowledge of the cooperating radar locations. Prior research has explored various synchronization strategies, including wired [13], GPS (GNSS)-based [22, 21, 3], and GPS-independent wireless approaches [2, 15, 23].
@@
In GPS-based approaches, an oscillator drives the digital subsystems, such as sampling-clock PLLs (ADC/DACs), carrier-generating PLLs (LO), and data clocks, and is disciplined by a GPS timing receiver, commonly referred to as a GPS-disciplined oscillator (GPSDO). These units provide stable timing and frequency references, with typical timing accuracies on the order of nanoseconds and fractional frequency stability on the order of $10^{-13}$$10^{-11}$ [3, 14], achieving long-term stability. The short-term stability is determined by the oscillator used and its phase-noise characteristics. However, GPS signals can be easily spoofed or jammed, require LOS visibility to satellites, and typically need a long acquisition time to achieve lock. Moreover, high-performance GPSDOs are expensive, although low-cost commercial off-the-shelf (COTS) options are available with moderate stability.
@@
To overcome the dependency on GPS, independent wireless synchronization approaches have been developed, which can be broadly classified into centralized [2, 23, 15] and decentralized [24] architectures. These are designed for general distributed wireless communication systems, with a focus on improving synchronization performance and demonstrated sub-hertz frequency accuracy and nanosecond-to-picosecond (ns–ps) time accuracy. In centralized systems, a leader–follower architecture is adopted, where the followers synchronize to the leader by exchanging frequency and time reference signals. Frequency synchronization techniques typically rely on the wireless transmission of an analog [1, 2, 15, 5] or digitally modulated reference from the leader. Time synchronization, on the other hand, is achieved through bi-directional waveform (BDW) exhanges [2, 23, 15, 4, 6, 18], which can be implemented using either timestamp-based or timestamp-free protocols. A key challenge for frequency synchronization is its susceptibility to multi-path and Doppler effects, while time synchronization schemes require symmetric propagation delays between the leader and followers. Furthermore, the achievable accuracy is strongly dependent on the SNR of the wireless link.
@@
Prior studies have investigated GPS-based synchronization for MSR from different perspectives. Coherent network radar operation using custom GPSDOs was demonstrated in [22]. The follow-up work [21] analyzed synchronization limitations and proposed LOS phase compensation for bistatic coherence. The work in [3] evaluated commercial GPSDOs for practical radar synchronization. Wireless adaptation of the white rabbit precision time protocol (WR-PTP) was studied in [19, 8], with a particular focus on synchronization. In contrast, this paper investigates the performance of a leader–follower MSR architecture employing GPS-independent wireless synchronization. A Bayesian Cramér–Rao lower bound (BCRLB) [20, 25, 26] framework is developed to study the impact of the residual synchronization offsets on target position and velocity estimation accuracy. In this work, frequency synchronization is achieved by extracting a reference clock from a two-tone waveform (TTW) broadcasted by leader to all followers [1, 2, 15, 5]. Time synchronization is accomplished using a timestamp-free bidirectional waveform exchange that estimates and corrects clock offsets while compensating for propagation delays [4, 6, 18]. Simulation results demonstrate the MSR performance under multiple configurations of synchronization-link channel and transmit parameters, and compared with the monostatic radar.
@@
The paper is structured as follows. First, the system model incorporating synchronization offsets is introduced. Next, the BCRLB framework is presented, where the synchronization offsets are treated as nuisance parameters with prior distributions. The subsequent section describes the MSR geometry and derives the corresponding CRLBs for target position and velocity. The procedures for frequency and time synchronization are then outlined. Finally, the simulation results are discussed, emphasizing the synchronization integrated MSR’s performance, followed by the conclusion and future works.
@@ II Signal Model with Synchronization Offsets
In this section, we present the signal model incorporating synchronization offsets along with the underlying assumptions. Fig. 1 illustrates a MSR architecture featuring inter-radar wireless links for synchronization and data exchange, as well as radar links between the radars and the target. The followers communicate wirelessly with the leader to achieve synchronization and to exchange Time of Arrival (TOA) and Frequency of Arrival (FOA) measurements. The radars performing measurements are assumed to be either mobile or fixed wireless nodes with known locations relative to the leader. The focus of this work is on the impact of synchronization errors, while the position uncertainty of the followers is neglected. The followers, considered as passive radars (e.g., mounted on drones or vehicular nodes), are geographically distributed around the leader, providing scalability in the number of nodes and adaptability to mobility based on operational requirements.
The transmitted signal from the leader radar is expressed as $s_{p}(t)=s(t)e^{j2\pi f_{c}t},$ where $s(t)$ denotes the complex baseband radar pulse with a duration of $T$ seconds, and $f_{c}$ is the carrier frequency. Assuming a LOS channel between the transmitter and target as well as receiver and target and neglecting synchronization offsets, the received signal at the $i$th follower after down-conversion can be written as
@@
In practical systems, synchronization is typically performed parallel or in advance to target detection, to make sure that the carrier frequency offsets (CFOs) and clock (time) offsets (TOs) are corrected, ensuring carrier and timing alignment among all cooperating radars. However, residual CFOs and TOs may still remain, resulting in frequency and timing mismatches across the radars. Incorporating these residual offsets into the ideal received signal and accounting for complex additive white Gaussian noise (AWGN), $n_{i}(t)\sim\mathcal{CN}(0,N_{0})$, the received signal at the $i$th radar is
@@
where $\delta t_{i}$ and $\delta f_{i}$ represent the residual TOs and CFOs, respectively and will be treated as nuisance parameters [9]. We combine the amplitude term $\sqrt{G_{i}}$ and the constant phasor $h_{i}=h_{i}^{r}+jh_{i}^{i}=e^{j(-2\pi(f_{i}+\delta f_{i})\tau_{i}+\theta_{i})}$ into a single complex channel coefficient $A_{i}\triangleq\sqrt{G_{i}}h_{i}$. The received signal can thus be expressed as
@@
The terms $G_{i}$ and $h_{i}$ are included in the effective SNR, and since $|h_{i}|^{2}=1$, the received SNR is given by $\gamma_{i}=G_{i}\frac{E_{s}}{N_{0}}$. The deterministic parameters of interest are the TOA, $\tau_{i}$, and the FOA, $f_{i}$. The residual TOs and CFOs are modeled as independent Gaussian random variables, $\delta t_{i}\sim\mathcal{N}(0,\sigma_{t_{i}}^{2})$ and $\delta f_{i}\sim\mathcal{N}(0,\sigma_{f_{i}}^{2})$, with variances determined by the performance of the employed wireless synchronization techniques. The deterministic mean of the received (noise-free) signal is then given by $\mu_{i}(t)\triangleq\mathbb{E}\{r_{i}(t)\}=A_{i}\,s\!\left(t-(\tau_{i}+\delta t_{i})\right)e^{j2\pi(f_{i}+\delta f_{i})t}$.
@@ III BCRLB for Delay $\tau$ and Doppler $f_{d}$
We present the lower bounds on the variances of delay and Doppler estimation in the presence of synchronization offsets. The BCLRB [20, 25, 26] provides a lower bound on the mean squared error (MSE) of a estimator when prior information about the unknown parameters is available. It extends the classical CRLB by incorporating information from both the likelihood and the parameters prior. Let $\boldsymbol{\theta}\in\mathbb{R}^{p}$ denote a random parameter vector of interest, and let $\boldsymbol{x}\in\mathbb{R}^{N}$ denote the observation vector obtained from a statistical model parameterized by $\boldsymbol{\theta}$, with $\hat{\boldsymbol{\theta}}(\boldsymbol{x})$ the estimator of $\boldsymbol{\theta}$. The BCRLB is defined as
$\mathbb{E}_{\boldsymbol{\theta},\boldsymbol{x}}\!\left[(\hat{\boldsymbol{\theta}}(\boldsymbol{x})-\boldsymbol{\theta})(\hat{\boldsymbol{\theta}}(\boldsymbol{x})-\boldsymbol{\theta})^{\top}\right]\succeq\mathbf{I}_{B}^{-1},$ (4)
where $\mathbf{I}_{B}\in\mathbb{R}^{p\times p}$ is the Bayesian information matrix (BIM), defined as
$\mathbf{I}_{B}\triangleq\mathbb{E}_{\boldsymbol{\theta},\boldsymbol{x}}\!\left[\nabla_{\boldsymbol{\theta}}\log p(\boldsymbol{x},\boldsymbol{\theta})\,\nabla_{\boldsymbol{\theta}}^{\top}\log p(\boldsymbol{x},\boldsymbol{\theta})\right].$ (5)
The $I_{B}$ can be decomposed into two components, $\mathbf{I}_{P}+\mathbf{I}_{N},$ where $\mathbf{I}_{P}$ represents the prior information term, given by
$\mathbf{I}_{P}\triangleq\mathbb{E}_{\boldsymbol{\theta}}\!\left[\nabla_{\boldsymbol{\theta}}\log p(\boldsymbol{\theta})\,\nabla_{\boldsymbol{\theta}}^{\top}\log p(\boldsymbol{\theta})\right],$ (6)
and $\mathbf{I}_{N}$ denotes the Fisher information matrix (FIM) averaged over the parameters prior,
$\mathbf{I}_{N}\triangleq\mathbb{E}_{\boldsymbol{\theta},\boldsymbol{x}}\!\left[\nabla_{\boldsymbol{\theta}}\log p(\boldsymbol{x}|\boldsymbol{\theta})\,\nabla_{\boldsymbol{\theta}}^{\top}\log p(\boldsymbol{x}|\boldsymbol{\theta})\right]=\mathbb{E}_{\boldsymbol{\theta}}\!\left[\mathbf{I}_{F}(\boldsymbol{\theta})\right],$ (7)
with FIM as $\mathbf{I}_{F}(\boldsymbol{\theta})\triangleq\mathbb{E}_{\boldsymbol{x}|\boldsymbol{\theta}}\!\left[\nabla_{\boldsymbol{\theta}}\log p(\boldsymbol{x}|\boldsymbol{\theta})\,\nabla_{\boldsymbol{\theta}}^{\top}\log p(\boldsymbol{x}|\boldsymbol{\theta})\right].$
@@ III-A1 Fisher Information
The FIM for the parameter vector $\boldsymbol{\theta}=[\tau_{i},f_{i},\delta t_{i},\delta f_{i}]$ is defined as
$\mathbf{I}_{F}^{i}(\boldsymbol{\theta})=\begin{bmatrix}I_{\tau_{i}\tau_{i}}&I_{\tau_{i}f_{i}}&I_{\tau_{i}\delta t_{i}}&I_{\tau_{i}\delta f_{i}}\\ I_{f_{i}\tau_{i}}&I_{f_{i}f_{i}}&I_{f_{i}\delta t_{i}}&I_{f_{i}\delta f_{i}}\\ I_{\delta t_{i}\tau_{i}}&I_{\delta t_{i}f_{i}}&I_{\delta t_{i}\delta t_{i}}&I_{\delta t_{i}\delta f_{i}}\\ I_{\delta f_{i}\tau_{i}}&I_{\delta f_{i}f_{i}}&I_{\delta f_{i}\delta t_{i}}&I_{\delta f_{i}\delta f_{i}}\end{bmatrix}$.
@@ III-A3 Full FIM with Nuisance Parameters
Since $\frac{\partial\mu_{i}}{\partial\tau_{i}}=\frac{\partial\mu_{i}}{\partial\delta t_{i}}$ and $\frac{\partial\mu_{i}}{\partial f_{i}}=\frac{\partial\mu_{i}}{\partial\delta f_{i}}$, the FIM can be written as $\mathbf{I}_{F}^{i}=\begin{bmatrix}\mathbf{A}^{i}&\mathbf{A}^{i}\\ \mathbf{A}^{i}&\mathbf{A}^{i}\end{bmatrix}$, where $\mathbf{A}^{i}$ denotes the FIM corresponding to the delay–Doppler parameters. In this work, the transmitted radar waveform is assumed to be a sequence of linear frequency modulated (LFM) pulses. The delay–Doppler FIM expressions derived in [7] are used directly to avoid re-derivation. The FIM is given by
$\mathbf{A}^{i}=2P\mathrm{SNR}_{i}\begin{bmatrix}\frac{1}{3}\pi^{2}f_{B}^{2}&-\frac{1}{6}\pi f_{B}T_{0}\\ -\frac{1}{6}\pi f_{B}T_{0}&\frac{1}{12}T_{0}^{2}\left(1+\left(\frac{T_{R}}{T_{0}}\right)^{2}(P^{2}-1)\right)\end{bmatrix}$ (9)
where $\mathrm{SNR}_{i}=\gamma_{i}\,\mathbf{a}(\phi)^{*}\mathbf{N}_{\Sigma}^{-1}\mathbf{a}(\phi)$ is the (post-processed [15]) SNR for a single pulse. Here, $P$ denotes the number of pulses, $T_{0}$ is the single-pulse duration, $f_{B}$ is the bandwidth of the chirp pulse, $T_{R}$ is the pulse repetition interval, $\mathbf{a}(\phi)$ is the array response vector, and $\mathbf{N}_{\Sigma}$ represents the spatial noise covariance matrix. Here, $\mathbf{I}_{N}^{i}=\mathbf{I}_{F}^{i}$, as $\mathbf{I}_{F}^{i}$ does not depend on the parameters of interest.